# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ II Multi-orbital TPSC
The spin independent TPSC3 Ansätze are defined as (fixing the external spin to $s_{1}=\uparrow$)
$\tilde{D}_{o_{1},o_{4}}$ $=D_{o_{1},o_{4}}\frac{\bar{\delta}_{o_{1},o_{4}}\braket{n^{\uparrow}_{o_{4}}n^{\uparrow}_{o_{1}}}+\braket{n^{\uparrow}_{o_{1}}n^{\downarrow}_{o_{4}}}}{\sum_{s_{4}}\braket{n_{o_{4}}^{s_{4}}}\braket{n_{o_{1}}^{\uparrow}}-\braket{n_{o_{1},o_{4}}^{\uparrow s_{4}}}\braket{n_{o_{4},o_{1}}^{s_{4}\uparrow}}}\;,$ (9)
$\tilde{C}_{o_{1},o_{4}}$ $=C_{o_{1},o_{4}}\frac{\bar{\delta}_{o_{1},o_{4}}\braket{n^{\uparrow}_{o_{4}}n^{\uparrow}_{o_{1}}}+\braket{n^{\uparrow\downarrow}_{o_{4}}n^{\downarrow\uparrow}_{o_{1}}}}{\sum_{s_{4}}\braket{n_{o_{4}}^{\uparrow,s_{4}}}\braket{n_{o_{1}}^{s_{4},\uparrow}}-\braket{n_{o_{1},o_{4}}^{s_{4}}}\braket{n_{o_{4},o_{1}}^{\uparrow}}}\;,$ (10)
$\tilde{P}_{o_{1},o_{4}}$ $=P_{o_{1},o_{4}}\frac{-\braket{n^{\uparrow\downarrow}_{o_{4}o_{1}}n^{\downarrow\uparrow}_{o_{4},o_{1}}}}{\braket{n_{o_{4}o_{1}}^{\uparrow,\downarrow}}\braket{n_{o_{4}o_{1}}^{\downarrow,\uparrow}}-\braket{n_{o_{4},o_{1}}^{\downarrow}}\braket{n_{o_{4},o_{1}}^{\uparrow}}}\;,$ (11)
while in the TPSC5 case the explicitly spin dependent Ansätze read
$\tilde{D}^{s_{1},s_{4}}_{o_{1},o_{4}}$ $=D_{o_{1},o_{4}}\bar{\delta}_{s_{1},s_{4}}\bar{\delta}_{o_{1},o_{4}}\frac{\braket{n^{s_{4}}_{o_{4}}n^{s_{1}}_{o_{1}}}}{\braket{n_{o_{4}}^{s_{4}}}\braket{n_{o_{1}}^{s_{1}}}-\braket{n_{o_{1},o_{4}}^{s_{1},s_{4}}}\braket{n_{o_{4},o_{1}}^{s_{4},s_{1}}}}\;,$ (12)
$\tilde{C}^{s_{1},s_{4}}_{o_{1},o_{4}}$ $=C_{o_{1},o_{4}}\bar{\delta}_{s_{1},s_{4}}\bar{\delta}_{o_{1},o_{4}}\frac{\braket{n^{s_{1},s_{4}}_{o_{4}}n^{s_{4},s_{1}}_{o_{1}}}}{\braket{n_{o_{4}}^{s_{1},s_{4}}}\braket{n_{o_{1}}^{s_{4},s_{1}}}-\braket{n_{o_{1},o_{4}}^{s_{4}}}\braket{n_{o_{4},o_{1}}^{s_{1}}}}\;,$ (13)
$\tilde{P}^{s_{1}}_{o_{1},o_{4}}$ $=P_{o_{1},o_{4}}\frac{-\braket{n^{s_{1},\bar{s}_{1}}_{o_{4}o_{1}}n^{\bar{s}_{1},s_{1}}_{o_{4},o_{1}}}}{\braket{n_{o_{4}o_{1}}^{s_{1},\bar{s}_{1}}}\braket{n_{o_{4}o_{1}}^{\bar{s}_{1},s_{1}}}-\braket{n_{o_{4},o_{1}}^{\bar{s}_{1}}}\braket{n_{o_{4},o_{1}}^{s_{1}}}}\;,$ (14)
@@
where we introduced the short hand notation $n_{o_{1},o_{2}}^{s_{1},s_{2}}=c^{\dagger}_{o_{2},s_{2}}(\tau,\bm{r})c_{o_{1},s_{1}}(\tau,\bm{r})$. Additionally the second index is dropped whenever two indices are identical. In principle, the two approaches should yield compatible results as long as the underlying assumptions of the approach are valid. Furthermore, it should be noted that if the initial model contains no inter-orbial hopping between orbitals which are interacting, the renormalizations of $\tilde{P}$ for both TPSC3 and TPSC5 as well as $\tilde{C}^{\uparrow,\downarrow}$ in TPSC5 are singular and thus the local and static limits cannot be captured in the standard fashion. Therefore, the question arises on how to renormalize these components in such a case. The first option is to stay at the level of plain Hartree-Fock (which does not renormalize these couplings leading to an overestimation of their contribution). The second option is to use the freedom of the Ansatz and to use the same rescaling as we use for the other components. In principle, as long as the trace consistency check (the local and static limit of Eq. (7)) is fulfilled, both variants are reasonable, however the former is expected to break down rapidly at stronger coupling, while the latter maintains the Kanamori-Brückner-like scaling of TPSC [78].
@@
As is usual for TPSC, the unknown local and static TPEVs are determined self-consistently. Specifically, the TPEVs are calculated from sum rules, for which one requires the susceptibility which we determine utilizing the Bethe-Salpeter equation. The interaction which is put into the Bethe-Salpeter equation is derived as the functional derivative from the Ansatz equation – thus closing the self-consistency. In order to obtain the vertex, see Eq. (64) in Appendix A, one reformulates the equation of motion into an explicit equation for the self-energy, as shown here for the case of TPSC5 (the equations for TPSC3 are obtained by using the symmetries between the different components):
$\Sigma_{o_{1},o_{3}}^{s_{1},s_{3}}$ $=\Big{(}-\sum_{s_{4},o_{4}}\tilde{D}_{o_{4},o_{1}}^{s_{4},s_{1}}n_{o_{4}}^{s_{4}}\delta_{s_{1},s_{3}}\delta_{o_{1},o_{3}}+\tilde{D}_{o_{3},o_{1}}^{s_{3},s_{1}}G_{o_{1},o_{3}}^{s_{1},s_{3}}+\sum_{o_{4}}\tilde{C}^{s_{1},s_{3}}_{o_{1},o_{4}}G_{o_{4},o_{4}}^{s_{1},s_{3}}\delta_{o_{1},o_{3}}$ (15)
$\qquad-\sum_{s_{4}}\tilde{C}^{s_{1},s_{4}}_{o_{1},o_{3}}G_{o_{1},o_{3}}^{s_{4},s_{4}}\delta_{s_{1},s_{3}}+\tilde{P}^{s_{1},\bar{s}_{1}}_{o_{3},o_{1}}G_{o_{3},o_{1}}^{s_{1},\bar{s}_{1}}\delta_{\bar{s}_{1},s_{3}}-\tilde{P}^{s_{1},\bar{s}_{1}}_{o_{3},o_{1}}G_{o_{3},o_{1}}^{\bar{s}_{1},\bar{s}_{1}}\delta_{s_{1},s_{3}}\Big{)}.$ (16)
@@
The irreducible vertex is defined as the functional derivative of the self-energy w.r.t the Green’s function
$\Gamma_{1,2,3,4}=\frac{\delta\Sigma_{1,2}}{\delta G_{4,3}}\,.$ (17)
@@
Therefore, the vertex contains functional derivatives of the Ansätze themselves which are a priori unknown. The way to circumvent this is to calculate the vertex in the Pauli matrix basis [78], transforming spin into $n=S^{0}$, $S^{x}=S^{1}$, $S^{y}=S^{2}$ and $S^{z}=S^{3}$. The transformation between the Pauli ($\Gamma^{S^{i^{\prime}}S^{j^{\prime}}}$) and diagrammatic ($\Gamma^{s_{1},s_{2},s_{3},s_{4}}$) spin space is defined by the Pauli matrices,
$\Gamma^{S^{i}S^{j}}=\frac{1}{2}\sum_{s_{1},s_{2},s_{3},s_{4}}\sigma^{i}_{s_{1},s_{3}}\Gamma^{s_{1},s_{2},s_{3},s_{4}}\sigma^{i}_{s_{2},s_{4}}.$ (18)
Thus the irreducible vertex in physical spin space in the crossed particle-hole channel is
$\Gamma^{S^{i}S^{j}}_{o_{1}o_{2}o_{3}o_{4}}=\frac{1}{2}\sum_{s_{1},s_{2},s_{3},s_{4}}\sigma^{i}_{s_{1},s_{3}}\frac{\delta\Sigma_{o_{1},o_{3}}^{s_{1},s_{3}}}{G_{o_{4},o_{2}}^{s_{4},s_{2}}}\sigma^{i}_{s_{2},s_{4}}\;.$ (19)
First, let us consider
$\Gamma^{S^{3}S^{3}}_{o_{1}o_{2}o_{3}o_{4}}$ $=\frac{1}{2}\Big{(}\frac{\delta\Sigma_{o_{1},o_{3}}^{\uparrow,\uparrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}+\frac{\delta\Sigma_{o_{1},o_{3}}^{\downarrow,\downarrow}}{\delta G_{o_{4},o_{2}}^{\downarrow,\downarrow}}-\frac{\delta\Sigma_{o_{1},o_{3}}^{\uparrow,\uparrow}}{\delta G_{o_{4},o_{2}}^{\downarrow,\downarrow}}-\frac{\delta\Sigma_{o_{1},o_{3}}^{\downarrow,\downarrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}\Big{)}.$ (20)
In an $SU(2)$ symmetric system, the Ansätze have the symmetry $X^{s_{1},s_{2}}=X^{\bar{s}_{1},\bar{s}_{2}}$ such that in each expression we can simultaneously flip all spins. Furthermore, there are two different classes of contributions, the ones where the functional derivative acts on an Ansatz and the ones in which it acts on a propagator. The former class occurs either as
$n_{o_{4}}^{s_{4}}{\delta}_{o_{1},o_{3}}\Big{(}\frac{\delta\tilde{D}_{o_{4},o_{1}}^{s_{4},\uparrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}-\frac{\delta\tilde{D}_{o_{4},o_{1}}^{s_{4},\downarrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}\Big{)}=n_{o_{4}}^{\uparrow}{\delta}_{o_{1},o_{3}}\Big{(}\frac{\delta\tilde{D}_{o_{4},o_{1}}^{\uparrow,\uparrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}-\frac{\delta\tilde{D}_{o_{4},o_{1}}^{\uparrow,\downarrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}\Big{)}+n_{o_{4}}^{\downarrow}{\delta}_{o_{1},o_{3}}\Big{(}\frac{\delta\tilde{D}_{o_{4},o_{1}}^{\downarrow,\uparrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}-\frac{\delta\tilde{D}_{o_{4},o_{1}}^{\downarrow,\downarrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}\Big{)},$ (21)
which is zero due to the symmetry of the vertex and the SU(2) symmetry of the occupation number, or as
$\sum_{o_{4}}G_{o_{4},o_{4}}^{\uparrow,\uparrow}\bar{\delta}_{o_{1},o_{3}}\Big{(}\frac{\delta\tilde{C}_{o_{1},o_{4}}^{\uparrow,\uparrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}-\frac{\delta\tilde{C}_{o_{1},o_{4}}^{\downarrow,\downarrow}}{\delta G_{o_{4},o_{2}}^{\uparrow,\uparrow}}\Big{)}\,,$ (22)
which cancels due to the spin symmetry of the Ansätze. Thus the only contributions we are left with are the ones where the derivative acts on the propagator, leaving us with
@@
$\Gamma^{S^{3}S^{3}}_{o_{1}o_{2}o_{3}o_{4}}$ $=-\delta_{o_{2},o_{4}}\delta_{o_{1},o_{3}}\tilde{D}^{\uparrow\uparrow}_{o_{4},o_{1}}+\delta_{o_{3},o_{2}}\delta_{o_{1},o_{4}}\tilde{D}^{\uparrow\uparrow}_{o_{4},o_{1}}+\delta_{o_{2},o_{4}}\delta_{o_{1},o_{3}}\tilde{C}^{\uparrow\uparrow}_{o_{1},o_{4}}-\delta_{o_{3},o_{2}}\delta_{o_{1},o_{4}}\tilde{C}^{\uparrow\uparrow}_{o_{1},o_{3}}$ (23)
$\quad+\delta_{o_{1},o_{3}}\delta_{o_{2},o_{4}}\tilde{D}^{\downarrow\uparrow}_{o_{4},o_{1}}+\tilde{C}^{\uparrow\downarrow}_{o_{1},o_{3}}\delta_{o_{3},o_{2}}\delta_{o_{1},o_{4}}+\tilde{P}_{o_{3},o_{1}}\delta_{o_{1},o_{2}}\delta_{o_{3},o_{4}}\,.$
@@
$\Gamma^{S^{3}S^{3}}_{o_{1}o_{2}o_{3}o_{4}}=\tilde{P}_{o_{3},o_{1}}\delta_{o_{1},o_{2}}\delta_{o_{3},o_{4}}+\tilde{C}_{o_{1},o_{4}}\delta_{o_{2},o_{4}}\delta_{o_{1},o_{3}}+\tilde{D}_{o_{4},o_{1}}\delta_{o_{3},o_{2}}\delta_{o_{1},o_{4}},$ (24)
which is again the standard form of the inter-orbital bilinear vertex. Notably, in TPSC5 we can define $X=\tilde{D}^{\uparrow\uparrow}-\tilde{C}^{\uparrow\uparrow}$, meaning that we technically do not have five but only four independent vertex components. While for the other diagonal spin components, the identical result is found, in the charge channel, no cancellation of the functional derivatives of the Ansätze exists, and hence this vertex cannot be calculated from these and we have to resort to a two-step procedure [78]. The procedure is as follows: first one obtains the TPEV from the self-consistent solution of the spin channel, while in a next step one fits the charge channel vertex such that the sum rules derived below are fulfilled.
@@
$\chi^{S^{3}S^{3}}_{o_{1},o_{3},o_{4},o_{2}}(\bm{q})=\chi^{0;S^{3}S^{3}}_{o_{1},o_{3},o_{4}^{\prime},o_{2}^{\prime}}(\bm{q})\left(\frac{1}{1-\chi^{0;S^{3}S^{3}}\Gamma^{S^{3}S^{3}}}\right)_{o_{4}^{\prime}o_{2}^{\prime},o_{4},o_{2}}(\bm{q})\,,$ (25)
where we shuffeled index $2$, $3$ and $4$ to obtain an equation in the form of a matrix-matrix product. The susceptibilities are then used to calculate the required TPEVs, which read
@@
These are then calculated by utilizing the orbitally-resolved sum rules of the susceptibility given by [93]
$\sum_{\bm{q}}\chi^{\alpha,\beta}_{o_{1}o_{3}o_{2}o_{4}}(\bm{q})=\chi^{\alpha,\beta}_{o_{1}o_{3}o_{2}o_{4}}=\braket{T_{\tau}\alpha_{13}\beta_{24}}-\braket{\alpha_{13}}\braket{\beta_{24}}.$ (27)
Here, $\alpha$ and $\beta$ are either spin operators or density operators. For spin operators in a $SU(2)$ symmetric system $\braket{\alpha_{13}}$ has to be zero, so that we can drop the single particle contribution in the following.
@@
$=2\braket{n_{o_{1}o_{3}}^{\uparrow}n_{o_{2}o_{4}}^{\uparrow}-n_{o_{1}o_{3}}^{\uparrow}n_{o_{2}o_{4}}^{\downarrow}}\,$ (29)
by differentiating cases with pairwise identical indices:
@@
where we utilized the Pauli principle in the special case where all indices are identical. However, these equations are not sufficient to fix all unknowns. To do so, we use the fact that the susceptibility is $SU(2)$ symmetric, i.e. $\chi^{S^{1}S^{1}}=\chi^{S^{2}S^{2}}=\chi^{S^{3}S^{3}}$. This allows us to obtain more sum rules by evaluating
@@
=±$=\pm\left(\braket{c^{\dagger}_{o_{3}\downarrow}c_{o_{1}\uparrow}c^{\dagger}_{o_{4}\downarrow}c_{o_{2}\uparrow}}+\braket{c^{\dagger}_{o_{3}\uparrow}c_{o_{1}\downarrow}c^{\dagger}_{o_{4}\uparrow}c_{o_{2}\downarrow}}\pm\braket{c^{\dagger}_{o_{3}\downarrow}c_{o_{1}\uparrow}c^{\dagger}_{o_{4}\uparrow}c_{o_{2}\downarrow}}\pm\braket{c^{\dagger}_{o_{3}\uparrow}c_{o_{1}\downarrow}c^{\dagger}_{o_{4}\downarrow}c_{o_{2}\uparrow}}\right),$ (35)

### v3

@@ 2 Multi-orbital TPSC
The single spin dependent TPSC3 Ansätze are defined as (fixing the external spin to $s_{1}=\,\uparrow$)
$\tilde{D}_{o_{4},o_{1}}$ $=D_{o_{4},o_{1}}\frac{\braket{n^{\uparrow}_{o_{4}}n^{\uparrow}_{o_{1}}}+\braket{n^{\uparrow}_{o_{1}}n^{\downarrow}_{o_{4}}}}{\sum_{s_{4}}\braket{n_{o_{4}}^{s_{4}}}\braket{n_{o_{1}}^{\uparrow}}-\braket{n_{o_{1},o_{4}}^{\uparrow,s_{4}}}\braket{n_{o_{4},o_{1}}^{s_{4},\uparrow}}}\;,$ (11)
$\tilde{C}_{o_{1},o_{3}}$ $=C_{o_{1},o_{3}}\frac{\braket{n^{\uparrow}_{o_{3}}n^{\uparrow}_{o_{1}}}+\braket{n^{\uparrow\downarrow}_{o_{3}}n^{\downarrow\uparrow}_{o_{1}}}}{\sum_{s_{4}}\braket{n_{o_{3}}^{\uparrow,s_{4}}}\braket{n_{o_{1}}^{s_{4},\uparrow}}-\braket{n_{o_{1},o_{3}}^{s_{4}}}\braket{n_{o_{3},o_{1}}^{\uparrow}}}\;,$ (12)
$\tilde{P}_{o_{1},o_{2}}$ $=P_{o_{1},o_{2}}\frac{\braket{n^{\uparrow\downarrow}_{o_{2}o_{1}}n^{\downarrow\uparrow}_{o_{2},o_{1}}}}{\braket{n_{o_{2},o_{1}}^{\uparrow,\downarrow}}\braket{n_{o_{2},o_{1}}^{\downarrow,\uparrow}}-\braket{n_{o_{2},o_{1}}^{\downarrow}}\braket{n_{o_{2},o_{1}}^{\uparrow}}}\;,$ (13)
where we dropped the dependence on the single spin as the $SU(2)$ symmetry of all quantities on the right hand side implies $X^{\uparrow}=X^{\downarrow}$ (where $X$ stands for an arbitrary ansatz). In the TPSC5 case the two-spin dependent Ansätze read
$\tilde{D}^{s_{4},s_{1}}_{o_{4},o_{1}}$ $=D_{o_{4},o_{1}}\frac{\braket{n^{s_{4}}_{o_{4}}n^{s_{1}}_{o_{1}}}}{\braket{n_{o_{4}}^{s_{4}}}\braket{n_{o_{1}}^{s_{1}}}-\braket{n_{o_{1},o_{4}}^{s_{1},s_{4}}}\braket{n_{o_{4},o_{1}}^{s_{4},s_{1}}}}\;,$ (14)
$\tilde{C}^{s_{1},s_{3}}_{o_{1},o_{3}}$ $=C_{o_{1},o_{3}}\frac{\braket{n^{s_{1},s_{3}}_{o_{3}}n^{s_{3},s_{1}}_{o_{1}}}}{\braket{n_{o_{3}}^{s_{1},s_{3}}}\braket{n_{o_{1}}^{s_{3},s_{1}}}-\braket{n_{o_{1},o_{3}}^{s_{3}}}\braket{n_{o_{3},o_{1}}^{s_{1}}}}\;,$ (15)
$\tilde{P}^{s_{1}}_{o_{2},o_{1}}$ $=P_{o_{2},o_{1}}\frac{\braket{n^{s_{1},\bar{s}_{1}}_{o_{2}o_{1}}n^{\bar{s}_{1},s_{1}}_{o_{2},o_{1}}}}{\braket{n_{o_{2},o_{1}}^{s_{1},\bar{s}_{1}}}\braket{n_{o_{2},o_{1}}^{\bar{s}_{1},s_{1}}}-\braket{n_{o_{2},o_{1}}^{\bar{s}_{1}}}\braket{n_{o_{2},o_{1}}^{s_{1}}}}\;.$ (16)
In these equations, we introduced a short hand notation which allows us to drop redundant arguments, i.e. $n_{o_{1},o_{2}}^{s_{1},s_{2}}=c^{\dagger}_{o_{2},s_{2}}(\tau,\boldsymbol{r})c_{o_{1},s_{1}}(\tau,\boldsymbol{r})$ and
@@
$n_{o_{1}}^{s_{1}}=n_{o_{1},o_{1}}^{s_{1},s_{1}}=c^{\dagger}_{o_{1},s_{1}}(\tau,\boldsymbol{r})c_{o_{1},s_{1}}(\tau,\boldsymbol{r})$. In principle, the two approaches should yield compatible results as long as the underlying assumptions of the approach are valid. Furthermore, it should be noted that if the initial model contains no inter-orbial hopping between orbitals which are interacting, the renormalizations of $\tilde{P}$ for both TPSC3 and TPSC5 as well as $\tilde{C}^{\uparrow,\downarrow}$ in TPSC5 are singular and thus the local and static limits cannot be captured in the standard fashion. Therefore, the question arises on how to renormalize these components in such a case. The first option is to stay at the level of plain Hartree-Fock which does not renormalize these couplings leading to an overestimation of their contribution. The second option is to use the freedom of the Ansatz and employ the same rescaling as we use for the other components. In principle, as long as the trace consistency check (equivalence of left and right hand side in Eq. (10)) is fulfilled, both variants are reasonable. However the former is expected to break down rapidly at stronger coupling, while the latter maintains the Kanamori-Brückner-like scaling [80]. We will maintain the first variant for easier comparability with earlier multi-orbital TPSC approaches [89, 86].
@@
As is usual for TPSC [80, 82, 86], the unknown local and static TPEVs are determined self-consistently. Specifically, the TPEVs are calculated from susceptibility sum-rules, for which one requires the susceptibility, which we determine using the Bethe-Salpeter equation. The interaction which is put into the Bethe-Salpeter equation is derived as the functional derivative from the Ansatz equation – thus closing the self-consistency. In order to obtain the two-particle interaction, see Eq. (82) in Appendix A, one reformulates the equation of motion into an explicit equation for the self-energy, as shown here for the case of TPSC5 (the equations for TPSC3 are obtained by using the symmetries between the different components of the Ansätze):
$\Sigma_{o_{1},o_{3}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3})$ $=\delta_{\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3}}\Big{(}\tilde{D}_{o_{4}^{\prime},o_{1}}^{s_{4}^{\prime},s_{1}}G_{o_{4}^{\prime},o_{4}^{\prime}}^{s_{4}^{\prime},s_{4}^{\prime}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})\delta_{s_{1},s_{3}}\delta_{o_{1},o_{3}}-\tilde{D}_{o_{3},o_{1}}^{s_{3},s_{1}}G_{o_{1},o_{3}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})$ (17)
$\qquad+\tilde{C}^{s_{1},s_{3}^{\prime}}_{o_{1},o_{3}}G_{o_{1},o_{3}}^{s_{3}^{\prime},s_{3}^{\prime}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})\delta_{s_{1},s_{3}}-\tilde{C}^{s_{1},s_{3}}_{o_{1},o_{3}^{\prime}}G_{o_{3}^{\prime},o_{3}^{\prime}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})\delta_{o_{1},o_{3}}$
$\qquad+\tilde{P}^{s_{1},\bar{s}_{1}}_{o_{3},o_{1}}G_{o_{3},o_{1}}^{\bar{s}_{1},\bar{s}_{1}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})\delta_{s_{1},s_{3}}-\tilde{P}^{s_{1},\bar{s}_{1}}_{o_{3},o_{1}}G_{o_{3},o_{1}}^{s_{1},\bar{s}_{1}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})\delta_{\bar{s}_{1},s_{3}}\Big{)}.$
It should be noted that the self-energy we arrive at is purely spin-diagonal (due to the Green’s function being spin diagonal) and $\Sigma^{\uparrow,\uparrow}=\Sigma^{\downarrow,\downarrow}$ holds due to the symmetry of the Hamiltonian which is inherited by the Ansatz.
@@
The irreducible vertex in the direct particle-hole channel is defined as the functional derivative of the self-energy w.r.t the Green’s function
$\Gamma_{o_{1}o_{2}o_{3}o_{4}}^{s_{1}s_{2}s_{3}s_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{4})=\frac{\delta\Sigma_{o_{1},o_{4}}^{s_{1},s_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{4})}{\delta G_{o_{3};o_{2}}^{s_{3},s_{2}}(\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{2})}\,.$ (18)
@@
Therefore, the vertex contains functional derivatives of the Ansätze themselves which are a priori unknown. The way to circumvent their appearance is to calculate the vertex in the Pauli matrix basis [80], transforming spin into $n=S^{0}$, $S^{x}=S^{1}$, $S^{y}=S^{2}$ and $S^{z}=S^{3}$. The transformation between the Pauli ($\Gamma^{S^{i^{\prime}}S^{j^{\prime}}}$) and diagrammatic ($\Gamma^{s_{1},s_{2},s_{3},s_{4}}$) spin space is defined by
$\Gamma^{S^{i}S^{j}}_{o_{1}o_{2}o_{3}o_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{4})=\frac{1}{2}\sum_{s_{1},s_{2},s_{3},s_{4}}\sigma^{i}_{s_{1},s_{4}}\Gamma^{s_{1},s_{2},s_{3},s_{4}}_{o_{1}o_{2}o_{3}o_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{4})\sigma^{j}_{\bar{s}_{2},\bar{s}_{3}}.$ (19)
Thus the irreducible vertex in physical spin space is
$\Gamma^{S^{i}S^{j}}_{o_{1}o_{2}o_{3}o_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{2},\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{4})=\frac{1}{2}\sum_{s_{1},s_{2},s_{3},s_{4}}\sigma^{i}_{s_{1},s_{4}}\frac{\delta\Sigma_{o_{1},o_{4}}^{s_{1},s_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{4})}{\delta G_{o_{3};o_{2}}^{s_{3},s_{2}}(\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{2})}\sigma^{j}_{\bar{s}_{2},\bar{s}_{3}}\;.$ (20)
A detailed derivation of the interaction from the self-energy is given in App. C, here we only show the final results of this calculation. For TPSC5 we find that the interaction is determined by
@@
$\Gamma^{S^{3}S^{3}}_{o_{1}o_{2}o_{3}o_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1})$ $=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\left(\tilde{D}_{o_{3},o_{1}}^{\uparrow,\downarrow}-\tilde{D}_{o_{3},o_{1}}^{\uparrow,\uparrow}\right)+\delta_{o_{1},o_{3}}\delta_{o_{2},o_{4}}\tilde{D}_{o_{4},o_{1}}^{\uparrow,\uparrow}$ (21)
$\;+\delta_{o_{1},o_{3}}\delta_{o_{2},o_{4}}(\tilde{C}^{\downarrow,\uparrow}_{o_{1},o_{4}}-\tilde{C}^{\uparrow,\uparrow}_{o_{1},o_{4}})+\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\tilde{C}^{\uparrow,\uparrow}_{o_{1},o_{3}}$
$\;+\delta_{o_{1},o_{2}}\delta_{o_{3},o_{4}}P_{o_{3},o_{1}}^{\uparrow,\downarrow}\,.$
@@
$\Gamma^{S^{3}S^{3}}_{o_{1}o_{2}o_{3}o_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1})=\tilde{P}_{o_{3},o_{1}}\delta_{o_{1},o_{2}}\delta_{o_{3},o_{4}}+\tilde{C}_{o_{1},o_{3}}\delta_{o_{2},o_{3}}\delta_{o_{1},o_{4}}+\tilde{D}_{o_{4},o_{1}}\delta_{o_{4},o_{2}}\delta_{o_{1},o_{3}},$ (22)
which is again the standard form of the inter-orbital bilinear vertex. Notably, for TPSC5 we observe that $\tilde{D}^{\uparrow\uparrow}$ and $\tilde{C}^{\uparrow\uparrow}$ do not appear independently in any of the equations, meaning that we technically do not have five but only four independent vertex components. We observe that both Ansätze fall into the required form of an $SU(2)$ symmetric interaction, justifying a posteriori that we kept both possibilities. While for the other diagonal spin components, the identical result is found, in the charge channel, no cancellation of the functional derivatives of the Ansätze exists. Hence this vertex cannot be calculated directly. Therefore, we have to resort to a two-step procedure [80]. The procedure is as follows: first one obtains the TPEV from the self-consistent solution of the spin channel, while in a next step one fits the charge channel vertex such that the sum rules derived below are fulfilled.
@@
$\chi^{S^{3}S^{3}}_{o_{1},o_{3},o_{2},o_{4}}(\boldsymbol{q})=\chi^{0;S^{3}S^{3}}_{o_{1},o_{3},o_{2^{\prime}},o_{4^{\prime}}}(\boldsymbol{q})\left(\frac{1}{1-\Gamma^{S^{3}S^{3}}\chi^{0;S^{3}S^{3}}}\right)_{o_{2^{\prime}}o_{4^{\prime}},o_{2},o_{4}}(\boldsymbol{q})\,,$ (23)
where we shuffeled index $2$, $3$ and $4$ compared to the Appendix C to obtain an equation in the form of a matrix-matrix product. The susceptibilities are then used to calculate the required TPEVs, which read
@@
These are calculated by utilizing the orbitally-resolved sum rules of the susceptibility given by [95]
$\sum_{\boldsymbol{q}}\chi^{\alpha,\beta}_{o_{1}o_{3}o_{2}o_{4}}(\boldsymbol{q})=\chi^{\alpha,\beta}_{o_{1}o_{3}o_{2}o_{4}}=\braket{T_{\tau}\alpha_{13}\beta_{24}}-\braket{\alpha_{13}}\braket{\beta_{24}}.$ (25)
Here, $\alpha$ and $\beta$ are either spin operators or density operators. For spin operators in an $SU(2)$ symmetric system $\braket{\alpha_{13}}$ has to be zero, so that we can drop the single particle contribution in the following.
@@
$=2\braket{n_{o_{1}o_{3}}^{\uparrow}n_{o_{2}o_{4}}^{\uparrow}-n_{o_{1}o_{3}}^{\uparrow}n_{o_{2}o_{4}}^{\downarrow}}\,.$ (27)
By differentiating cases with pairwise identical indices we arrive at the following set of equations
@@
where we utilized the Pauli principle in the special case where all indices are identical. However, these equations are not sufficient to fix all unknowns. To do so, we use the fact that the susceptibility is $SU(2)$ symmetric, i.e. $\chi^{S^{1}S^{1}}=\chi^{S^{2}S^{2}}=\chi^{S^{3}S^{3}}$. This allows us to obtain more sum rules by evaluating the expression for the other components (note that $S^{1/2}$ means $S^{1}$ for the upper sign and $S^{2}$ for the lower sign):
@@
=±$=\pm\left(\braket{c^{\dagger}_{o_{3}\downarrow}c_{o_{1}\uparrow}c^{\dagger}_{o_{4}\downarrow}c_{o_{2}\uparrow}}+\braket{c^{\dagger}_{o_{3}\uparrow}c_{o_{1}\downarrow}c^{\dagger}_{o_{4}\uparrow}c_{o_{2}\downarrow}}\right.$ (33)
±$\;\quad\pm\left.\braket{c^{\dagger}_{o_{3}\downarrow}c_{o_{1}\uparrow}c^{\dagger}_{o_{4}\uparrow}c_{o_{2}\downarrow}}\pm\braket{c^{\dagger}_{o_{3}\uparrow}c_{o_{1}\downarrow}c^{\dagger}_{o_{4}\downarrow}c_{o_{2}\uparrow}}\right),$